# ZoDi: Zero-Shot Domain Adaptation with Diffusion-Based Image Transfer

### v1

@@ Abstract
Deep learning models achieve high accuracy in segmentation tasks among others, yet domain shift often degrades the models’ performance, which can be critical in real-world scenarios where no target images are available. This paper proposes a zero-shot domain adaptation method based on diffusion models, called ZoDi, which is two-fold by the design: zero-shot image transfer and model adaptation. First, we utilize an off-the-shelf diffusion model to synthesize target-like images by transferring the domain of source images to the target domain. In this we specifically try to maintain the layout and content by utilising layout-to-image diffusion models with stochastic inversion. Secondly, we train the model using both source images and synthesized images with the original segmentation maps while maximizing the feature similarity of images from the two domains to learn domain-robust representations. Through experiments we show benefits of ZoDi in the task of image segmentation over state-of-the-art methods. It is also more applicable than existing CLIP-based methods because it assumes no specific backbone or models, and it enables to estimate the model’s performance without target images by inspecting generated images. Our implementation will be publicly available.
@@
Deep neural networks have greatly advanced the performance of various recognition tasks in computer vision [35, 38, 12, 16]. The recognition models typically perform well when the data distributions are consistent between the domains of training and testing set. Those come, however, with a downside that the performance may significantly drop when tested on out-of-distribution data [30, 43, 42], which can be critical in real-world applications.
@@
In this paper, we are also concerned ourselves with zero-shot domain adaptation applicable to general settings, but without restrictions or uncertainty of results. In particular, we propose a method called Zero-shot domain adaptation with Diffusion-Based Image Transfer (ZoDi) while focusing on the task of segmentation. As illustrated in Fig. 1, it comprises zero-shot image transfer and model adaptation. Our basic strategy is to leverage powerful diffusion models [14, 37] for zero-shot image transfer of source images to the target domain. That is, we first transfer images to the target domain using diffusion models although it involves a non-trivial question of how to keep the layout and content. We then train our model for segmentation while learning domain-robust representations by maximizing the similarity of image features from the two domains.
@@ 4.2 Main Results
For all the experiments of segmentation we report the mean Intersection over Union (mIoU) results by taking the mean of three independent trials on different random seeds. We discuss the results by our method, ZoDi, while comparing them against those by two state-of-the-art baselines: PØDA [8] and DATUM [1] (PØDA is a zero-shot domain adaptation method and DATUM is a one-shot, see Sec. 2 for more details). We also make a comparison to DAFormer [15] for a strong baseline of unsupervised domain adaptation where full images from the target domain are accessible. This work uses simple prompts to describe the target domains, e.g. “driving domain”. We also conduct source-only training with the same conditions for comparing ZoDi with the models trained only with source images. The results are summarised in Tab. 1.
@@
DAFormer [15] | 400 | ✓ | 30.3 | 44.5 | 41.0 | 58.2 | 47.9 |
DATUM+DAFormer [1] | one | ✓ | 19.9 | 43.0 | 39.0 | 60.3 | 44.7 |
PØDA [8] | zero | ✗ | 25.0 | 43.9 | 42.3 | 49.0 | 41.1 |
ZoDi (Ours) | zero | ✓ | 23.0 | 45.6 | 47.0 | 56.1 | 40.5 |
@@
We use the same architecture, prompts, data augmentation, and optimizer as in PØDA for fair comparisons. ZoDi outperforms PØDA in three settings by large margins (45.6 vs. 43.9 in clearsnow, 47.0 vs. 42.3 in clearrain, and 56.1 vs. 49.0 in clearfog). In the other two settings, daynight and realgame, PØDA performs better (23.0 vs. 25.0 in daynight and 40.5 vs. 41.1 in realgame). However, ZoDi performs better than the models trained only by source data (23.0 vs. 20.7 and 40.5 vs. 40.1, respectively), and the same is consistently true in all the settings. These results largely reflect the effectiveness of ZoDi.
When it comes to the settings where ZoDi does not outperform PØDA (daynight and realgame), it could be attributed to the difficulties in drastically changing the domains. As can be seen in Fig. 3, it can be challenging for ZoDi to make severe changes. For example, in daynight setting, ZoDi appears to find it hard to darken images to a large extent. In realgame setting, it struggles with the discrepancy between target images (as seen in Fig. 4) and generated images.
@@ 4.2.1 Qualitative Results of Zero-Shot Image Transfer:
We train DATUM with one random image from the target domain. As shown in Tab. 1, ZoDi outperforms DATUM with DAFormer in three settings (23.0 vs. 19.9 in daynight, 45.8 vs. 43.0 in clearsnow, and 47.0 vs. 39.0 in clearrain), while it is the contrary in other two settings (56.1 vs. 60.3 in clearfog and 40.5 vs. 44.7 in realgame). Notice that DATUM harms the original performance (20.7 to 19.9 in daynight and 39.4 to 39.0 in clearrain). Here, it is also worth noting that our ZoDi does not require any images from the target domain, nor does it rely on other complicated, unsupervised domain adaptation methods.
@@ 4.3 Ablation Studies
We compare our diffusion-based image transfer method with SDEdit [24] and InST [51]. SDEdit is equivalent to the method of our image transfer but without ControlNet and stochastic inversion, and InST is likewise without ControlNet. Fig. 5 shows the generated images by three methods including ours. Among those the qualitative results illustrate that our transfer method best maintains the original layout and contents. In our quantitative evaluation, as shown in Tab. 2, we replace our method with SDEdit or InST while keeping the model-level adaptation stage unchanged. Our image transfer method in ZoDi outperforms the other alternatives.
@@
Zero-shot image transfer | SDEdit [24] | 18.1 | 36.0 | 35.4 | 46.4 | 37.2 |
InST [51] | 19.6 | 34.8 | 37.6 | 51.1 | 37.6 | |
Similarity loss | w/o | 22.5 | 44.0 | 46.0 | 56.4 | 40.4 |
Full-version | ZoDi (Ours) | 23.0 | 45.6 | 47.0 | 56.1 | 40.5 |
@@ 4.4 Failure Cases
We study some failure cases of image transfer in ZoDi because the quality of generated images is of importance. As in Fig. 6, our method sometimes make errors such that it could generate other different objects, fail to generate objects correctly, not generate objects, or fail to maintain the original content of the images. In addition, in the daynight setting as in Fig. 6(a), it tends to fail to darken the sky as much as in the real target images (see top-left in Fig. 4).
@@ Appendix 0.A Hyper-parameter Strength
The generated images based on different strengths are shown in Fig. A. The figure illustrates that a greater value of strength makes bigger changes although it could result in losing some original content. We set the strength as for the fog and game domain and for the other domains: night, snow, and rain. We do not use images from the target domain to explore the hyper-parameters because it is possible by inspecting generated images.

### v2

@@ Abstract
Deep learning models achieve high accuracy in segmentation tasks among others, yet domain shift often degrades the models’ performance, which can be critical in real-world scenarios where no target images are available. This paper proposes a zero-shot domain adaptation method based on diffusion models, called ZoDi, which is two-fold by the design: zero-shot image transfer and model adaptation. First, we utilize an off-the-shelf diffusion model to synthesize target-like images by transferring the domain of source images to the target domain. In this we specifically try to maintain the layout and content by utilising layout-to-image diffusion models with stochastic inversion. Secondly, we train the model using both source images and synthesized images with the original segmentation maps while maximizing the feature similarity of images from the two domains to learn domain-robust representations. Through experiments we show benefits of ZoDi in the task of image segmentation over state-of-the-art methods. It is also more applicable than existing CLIP-based methods because it assumes no specific backbone or models, and it enables to estimate the model’s performance without target images by inspecting generated images. Our implementation will be publicly available at https://github.com/azuma164/ZoDi.
@@
Deep neural networks have greatly advanced the performance of various recognition tasks in computer vision [36, 39, 12, 16]. The recognition models typically perform well when the data distributions are consistent between the domains of training and testing set. Those come, however, with a downside that the performance may significantly drop when tested on out-of-distribution data [19, 50, 31, 44, 43], which can be critical in real-world applications.
@@
In this paper, we are also concerned ourselves with zero-shot domain adaptation applicable to general settings, but without restrictions or uncertainty of results. Inspired by the strong zero-shot abilities of diffusion models, we propose a method called Zero-shot domain adaptation with Diffusion-Based Image Transfer (ZoDi) while focusing on the task of segmentation. As illustrated in Fig. 1, it comprises zero-shot image transfer and model adaptation. Our basic strategy is to leverage powerful diffusion models [14, 38] for zero-shot image transfer of source images to the target domain. That is, we first transfer images to the target domain using diffusion models although it involves a non-trivial question of how to keep the layout and content. We then train our model for segmentation while learning domain-robust representations by maximizing the similarity of image features from the two domains.
@@ 4.2 Main Results
For all the experiments of segmentation, we report the mean Intersection over Union (mIoU) results by taking the mean of three independent trials on different random seeds. We discuss the results by our method, ZoDi, while comparing them with two state-of-the-art baselines: PØDA [8] and DATUM [1] (PØDA is a zero-shot domain adaptation method with CLIP and DATUM is a one-shot with diffusion models, see Sec. 2 for more details). We also make a comparison to DAFormer [15] for a strong baseline of unsupervised domain adaptation where full images from the target domain are accessible. This work uses simple prompts to describe the target domains, e.g. “driving domain”. We also conduct source-only training with the same conditions for comparing ZoDi with the models trained only with source images. The results are summarised in Tab. 1.
@@
| DAFormer [15] | 400 | ✓ | 30.3† | 44.5† | 41.0† | 58.2† | 47.9† |
| DATUM+DAFormer [1] | one | ✓ | 19.9† | 43.0† | 39.0† | 60.3† | 44.7† |
| PØDA [8] | zero | ✗ | 25.0 | 43.9 | 42.3 | 49.0† | 41.1 |
| ZoDi (Ours) | zero | ✓ | 24.7 | 45.6 | 47.0 | 56.1 | 40.5 |
@@
We use the same architecture, prompts, data augmentation, and optimizer as in PØDA for fair comparisons. ZoDi outperforms PØDA in three settings by large margins: 45.6 vs. 43.9 in clearsnow, 47.0 vs. 42.3 in clearrain, and 56.1 vs. 49.0 in clearfog. In the other two settings, daynight and realgame, PØDA shows a slight edge over ZoDi (24.7 vs. 25.0 and 40.5 vs. 41.1, respectively). However, it is important to note that the performance gap in these scenarios is relatively small. Despite PØDA’s marginally better results in these cases, ZoDi still surpasses models trained solely on source data (24.7 vs. 20.7 and 40.5 vs. 40.1, respectively). This trend holds consistently across all settings, highlighting ZoDi’s robust performance even in cases where it is slightly outperformed by PØDA.
When it comes to the settings where ZoDi does not outperform PØDA (daynight and realgame), it could be attributed to the difficulties in drastically changing the domains. As can be seen in Fig. 3, it can be challenging for ZoDi to make severe changes. For example, in daynight setting, ZoDi appears to find it hard to darken images to a large extent while keeping the image content. In realgame setting, it struggles with the discrepancy between target images (as seen in Fig. 4) and generated images.
@@ 4.2.1 Qualitative Results of Zero-Shot Image Transfer:
We train DATUM with one random image from the target domain. As shown in Tab. 1, ZoDi outperforms DATUM with DAFormer in three settings (24.7 vs. 19.9 in daynight, 45.8 vs. 43.0 in clearsnow, and 47.0 vs. 39.0 in clearrain), while it is the contrary in other two settings (56.1 vs. 60.3 in clearfog and 40.5 vs. 44.7 in realgame). Notice that DATUM harms the original performance (20.7 to 19.9 in daynight and 39.4 to 39.0 in clearrain). Here, it is also worth noting that our ZoDi does not require any images from the target domain, nor does it rely on other complicated, unsupervised domain adaptation methods.
@@ 4.3 Ablation Studies
We compare our diffusion-based image transfer method with SDEdit [25], InST [53], and ours without SI (Stochastic Inversion). SDEdit is equivalent to the method of our image transfer but without ControlNet and stochastic inversion, and InST is likewise without ControlNet. Fig. 5 shows the generated images by four methods including ours. Among those the qualitative results illustrate that our transfer method best maintains the original layout and contents. In our quantitative evaluation, as shown in Tab. 2, we replace our method with SDEdit, InST, or ours without SI while keeping the model-level adaptation stage unchanged. Our image transfer method in ZoDi outperforms the other alternatives. We also compare with other image transfer methods (Prompt-to-Prompt [13] and InstructPix2Pix) in supplementary.
@@
| Image transfer | SDEdit [25] | 18.2 | 36.0 | 35.4 | 46.4 | 37.2 |
| InST [53] | 16.3 | 34.8 | 37.6 | 51.1 | 37.6 | |
| w/o SI [53] | 24.3 | 43.1 | 45.0 | 53.9 | 39.5 | |
| Similarity loss | w/o | 24.5 | 44.0 | 46.0 | 56.4 | 40.4 |
| Full-version | ZoDi (Ours) | 24.7 | 45.6 | 47.0 | 56.1 | 40.5 |
@@ 4.4 Failure Cases
We study some failure cases of image transfer in ZoDi because the quality of generated images is of importance. As in Fig. 6, our method sometimes make errors such that it could generate other different objects, fail to generate objects correctly, not generate objects, or fail to maintain the original content of the images. In addition, in the daynight setting as in Fig. 6(b), it tends to fail to keep the original image content.
@@ Appendix 0.A Hyper-parameter Strength
The generated images based on different strengths are shown in Fig. A. The figure illustrates that a greater value of strength makes bigger changes although it could result in losing some original content. We set the strength as for the fog and game domain, for the snow, and rain domain, and for the night domain. We do not use images from the target domain to explore the hyper-parameters because it is possible by inspecting generated images.
Appendix 0.B Comparison with other image transfer methods
| CLIP | |||
| Domain | Method | Similarity | mIoU (%) |
| DayNight | Source-only | 20.5 | 8.2 |
| Prompt-to-Prompt [13] | 27.4 | 9.5 | |
| InstructPix2Pix [2] | 28.4 | 8.2 | |
| ZoDi (Ours) | 30.0 | 9.1 | |
| ClearSnow | Source-only | 20.2 | 14.7 |
| Prompt-to-Prompt [13] | 26.1 | 16.5 | |
| InstructPix2Pix [2] | 25.9 | 16.6 | |
| ZoDi (Ours) | 28.2 | 18.1 | |
| ClearRain | Source-only | 22.0 | 16.6 |
| Prompt-to-Prompt [13] | 27.7 | 18.2 | |
| InstructPix2Pix [2] | 28.5 | 19.2 | |
| ZoDi (Ours) | 29.7 | 18.9 | |
| ClearFog | Source-only | 22.0 | 16.4 |
| Prompt-to-Prompt [13] | 30.5 | 15.7 | |
| InstructPix2Pix [2] | 30.1 | 16.1 | |
| ZoDi (Ours) | 28.8 | 19.4 |
We also compare our image transfer methods with Prompt-to-Prompt [13] and InstructPix2Pix [2]. Our approach differs from others by focusing on image generation to train segmentation models. Unlike other methods, our approach generates images based on segmentation maps. Therefore, our method is more accurate against the segmentation maps (Tab. B). Especially, the layout of images collapses in the fog domain in other methods, while ZoDi succeeds in keeping the layout.
Given that Prompt-to-Prompt and InstructPix2Pix each take about 3 minutes to generate an image on an NVIDIA A100, generating all images corresponding to Cityscapes [6] would be prohibitively time-consuming. Therefore, we generate 100 images for each domain and each image transfer method. In Prompt-to-Prompt, we use “driving” as the source prompt and “driving domain” as the target prompt, the same as the prompt we used in Sec. 4.1. domain is a description of the target domain. In InstructPix2Pix, we use “make it domain” as the instruction prompt.
First, we evaluate the image transfer methods using CLIP similarity [34], which measures the similarity between the image and text prompt. The table A shows that our method best aligns with the text prompt on average. Second, we reported the mIoU score achieved by models trained using the original 100 images and the generated 100 images. ZoDi shows consistent improvement and outperforms other methods on average.
Our method not only excels in aligning generated images with textual prompts but also enhances the performance of segmentation models. This demonstrates our approach’s effectiveness in training segmentation models in domain adaptation settings.